# Quantum Hermitian conjugate and encoding unnormalized matrices

### v2

@@ III Matrix multiplication
$\hat{b}_{0}=b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1},$ (31)
$\hat{b}_{1}=b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0},$
$\hat{a}_{jk0}=\sum_{l=0}^{N-1}(a^{(1)}_{jl0}a^{(2)}_{lk0}-a^{(1)}_{jl1}a^{(2)}_{lk1}),$
$\hat{a}_{jk1}=\sum_{l=0}^{N-1}(a^{(1)}_{jl0}a^{(2)}_{lk1}+a^{(1)}_{jl1}a^{(2)}_{lk0}),$
so that $\hat{A}=A^{(1)}A^{(2)}=\{\hat{a}_{jk}:j,k=0,\dots,N-1\}$, $\hat{a}_{jk}=\hat{a}_{jk0}+i\hat{a}_{jk1}$, $\hat{b}=\hat{b}_{0}+i\hat{b}_{1}$, where $\hat{a}_{jkm}$, $\hat{b}_{m}$, $m=0,1$, are real numbers. We note that the first term in the final expression for $|\Phi_{5}\rangle$ plays the same role as the first term in the expression for $|\Phi_{0}\rangle$ in (16). It weakens the constraint on the normalization of the elements of the resulting matrix $\hat{A}$.
Figure 3: (a) The circuit for the matrix multiplication algorithm; $|\Phi_{out}\rangle$ is formed by subsystems $R_{1}$, $C_{2}$, $K_{1}$ and $M_{1}$; $X\equiv\sigma^{(x)}$, $Z\equiv\sigma^{(z)}$; we omit subscripts in notations $W^{(j)}$ for the brevity; operator $W^{(2)}$ is a novelty in comparison with earlier algorithm in ZBQKW_arXive2024 ; (b) Notation for multi-qubit CNOT.
III.1 Manipulations with input matrices
In this section we consider several manipulations with the input matrices $A^{(1)}$ and $A^{(2)}$. Instead of multiplication $A^{(1)}A^{(2)}$ we can perform multiplications $(A^{(1)})^{\dagger}A^{(2)}$, $A^{(1)}(A^{(2)})^{\dagger}$, $(A^{(1)})^{\dagger}(A^{(2)})^{\dagger}$, $(A^{(1)}A^{(2)})^{\dagger}=(A^{(2)})^{\dagger}(A^{(1)})^{\dagger}$ and change order of the multiplier. With this purpose we introduce the following operators.
To organize the Hermitian conjugate of $A^{(i)}$, $i=1,2$, we use operator (5) introduced in Sec.II.1:
@@ III.1 Manipulations with input matrices
To change position of matrices we apply the operator
$Q^{(3)}_{R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}}=SWAP_{R^{(1)}C^{(1)}}SWAP_{R^{(2)}C^{(2)}}SWAP_{M^{(1)}M^{(2)}}.$ (33)
@@
The circuit implementing these operators is shown in Fig.4. For instance, to calculate the product $A^{(1)}(A^{(2)})^{\dagger}$ we apply the operator $Q^{(2)}_{R^{(2)}C^{(2)}M^{(2)}}$ to the state $|\Phi_{0}\rangle$. To calculate $(A^{(1)}A^{(2)})^{\dagger})$ we apply the operators $Q^{(3)}_{R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}}$, $Q^{(2)}_{R^{(2)}C^{(2)}M^{(2)}}$, $Q^{(1)}_{R^{(1)}C^{(1)}M^{(1)}}$ to the state $|\Phi_{0}\rangle$. Application of operators $Q^{(i)}$, $i=1,2,3$, can be controlled by additional qubits $q^{(i)}$, i.e., by the following controlled operators:
$\hat{Q}^{(i)}_{S^{(i)}}=|1\rangle_{q^{(i)}}\,{}_{q^{(i)}}\langle 1|\otimes Q^{(i)}_{S^{(i)}}+|0\rangle_{q^{(i)}}\,{}_{q^{(i)}}\langle 0|\otimes I_{S^{(i)}},$ (34)
@@
where $S^{(i)}=R^{(i)}C^{(i)}M^{(i)}$, $i=1,2$, $S^{(3)}=R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}$ and $I_{S^{(i)}}$ is the identity operator applied to $S^{(i)}$. The state $|\varphi\rangle_{q}=|n_{1}\rangle_{q^{(1)}}|n_{2}\rangle_{q^{(2)}}|n_{3}\rangle_{q^{(3)}}$, $n_{i}=0,1$ ($i=1,2,3$) of the controlling qubits is not changed after running the algorithm.
@@
Of cause calculating $(A^{(1)}A^{(2)})^{\dagger}$ can be dome by performing Hermitian conjugation after multiplication $A^{(1)}A^{(2)}$, this is another option.
@@
Figure 4: Controlled manipulations with input matrices $A^{(1)}$ and $A^{(2)}$.
III.2 Measuring normalization $G$ and discussion
@@
Although the controlled measurement allows to resolve the problem of exponential decrease of the success probability with the number of qubits encoding considered matrices we lose the information about normalization $G$. However, this information might required in obtaining the final result. Below we explore the method of indirect measuring the normalization $G$.
@@
Let $S_{1}=|b^{(1)}b^{(2)}|^{2}$, this value is known due to the matrix preparation, see eq.(11). If, along with $S_{1}$, we know $S_{1}/G^{2}=\tilde{S}_{1}$, then $G=\sqrt{S_{1}/\tilde{S}_{1}}$. To find $\tilde{S}_{1}$, we refer to the state $|\Phi_{5}\rangle$ (30) and measure the state of the subsystem $K_{1}$ with the output 0. It is easy to understand that the probability of this output equals $\tilde{S}_{1}$. and can be found in result of multiple runs of the algorithm.
It is important that both introduced extensions do not decline the parameters of the multiplication circuit such as depth and space. Fig.3 shows that the depth is determined by the operator $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}$, while manipulations in Sec.III.1, see Fig.4, increase this depth only by insignificant constant factor. Thus, the estimation of the depth remains $O(\log N)$ similar to the estimation in ZBQKW_arXive2024 . Regarding the space, the circuit in Fig.3 includes only five additional one-qubit subsystems ($K^{(i)}$, $M^{(i)}$, $i=1,2$ , and $\tilde{B}$) in comparison with the circuit in ZBQKW_arXive2024 with one of them (the ancilla $\tilde{B}$) is needed for organization of controlled measurement. Therefore, the space estimation remains $O(\log N)$ as well.
@@
We have to note that the above measurement destroys the part of superposition quantum state in which the result of matrix multiplication is stored. Therefore, to get the full information, i.e., normalization $G$ and matrix product, we have to use the set of runs to probabilistically determine $G$ and then run the algorithm one more time to end up with the state $|\Phi_{6}\rangle$ encoding the product $A^{(1)}A^{(2)}$. However, if we don’t need $G$, then the single run is enough.
IV Conclusion
@@
In this paper we continue studying matrix-manipulation algorithms that are proposed in ZQKW_2024 ; ZBQKW_arXive2024 and based on the encoding the matrix elements into the pure superposition state of a quantum system. Although those algorithms work with complex numbers in general, the complex conjugation (in particular, Hermitian conjugation) can not be realized in frames of those settings. This disadvantage is removed by the first our extension which consists in separating the real and imaginary parts of the complex matrix under consideration by encoding these two parts into different probability amplitudes of a pure superposition state. Such extension allows constructing effective subroutine for operations typical for complex numbers, which are complex and Hermitian conjugations. However, this separation of real and imaginary parts requires proper modifications of the quantum algorithms for matrix addition, multiplication, determinant calculation, matrix inversion and solving linear systems that are described in ZQKW_2024 ; ZBQKW_arXive2024 . In our paper, we present such modification for the multiplication algorithm and demonstrate manipulation with the input matrices using Hermitian conjugation subroutine.
@@
Another disadvantage of the algorithms in Refs.ZQKW_2024 ; ZBQKW_arXive2024 is normalization constraint which is generated by the normalization condition for the superposition pure quantum state. This requirement is weakened by the second our extension which adds an additional term in the superposition state encoding a particular matrix. This additional term leads to replacing the normalization equality (2) with inequality (7) that imposes only the upper boundary on the sum of the squared absolute values of the matrix elements.
@@
Of course, each of these modifications requires involving an extra qubit. Therefore, along with two subsystems $R^{(i)}$ and $C^{(i)}$, $i=1,2$, enumerating rows and columns of the matrices $A^{(i)}$, $i=1,2$, the $i$th matrix is associated with the one-qubit subsystem $M^{(i)}$, responsible for separate encoding of the real and imaginary parts of complex matrix elements, and the one-qubit subsystem $K^{(i)}$ allowing to introduce an extra term into the superposition state encoding the matrix $A^{(i)}$, namely the presence of this extra term weakens the normalization constraint. Thus, only four additional qubits appear in the modified multiplication algorithm. We note that the second one-qubit ancilae $\tilde{B}$ is needed for organization the controlled measurement which removes the problem of small success probability that appears in the case of direct uncontrolled measurement of the ancilla stat.
@@
We emphasize that both extensions do not significantly decline the depth and space of the multiplication algorithm. Estimations for both of them remain $O(\log N)$, which is similar to the appropriate characteristics of the multiplication algorithm in ZBQKW_arXive2024 .
Similar modifications can be included in other matrix-manipulation algorithms considered in Refs.ZQKW_2024 ; ZBQKW_arXive2024 which is postponed for future study.

### v4

@@ V Matrix multiplication
$\hat{b}_{0}=b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1},\;\;\hat{b}_{1}=b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0}$, (70)
$\hat{a}_{jk0}=\sum_{l=0}^{N-1}(a^{(1)}_{jl0}a^{(2)}_{lk0}-a^{(1)}_{jl1}a^{(2)}_{lk1}),\;\;\hat{a}_{jk1}=\sum_{l=0}^{N-1}(a^{(1)}_{jl0}a^{(2)}_{lk1}+a^{(1)}_{jl1}a^{(2)}_{lk0})$,
so that $\hat{A}=A^{(1)}A^{(2)}=\{\hat{a}_{jk}:j,k=0,\dots,N-1\}$, $\hat{a}_{jk}=\hat{a}_{jk0}+i\hat{a}_{jk1}$, $\hat{b}=\hat{b}_{0}+i\hat{b}_{1}$, where $\hat{a}_{jkm}$, $\hat{b}_{m}$, $m=0,1$, are real numbers. We note that the first term in the state $|\Psi_{out}\rangle$ given in Eq.(69) plays the same role as the first term in the expressions for $|\Psi^{(l)}\rangle$, given in Eq.(54). It weakens the constraint on the normalization of the elements of the resulting matrix $\hat{A}$.
The weak point is that the probability of measuring the desired state $|1\rangle_{B_{2}}$ is small, $p=G^{2}/2^{n+1}$, and exponentially decreases with an increase in $n$. Therefore, to obtain the desired result $|1\rangle_{B_{2}}$ we have to run the algorithm $O(2^{n+1}/G^{2})$ times. However, we can replace the usual measurement by the controlled one, i.e., let us perform the controlled measurement of the ancilla qubit $B_{2}$ via the operator (see also Eq.(14) in Sec.II.2.2)
$W^{(4)}_{B}=|1\rangle_{B_{1}}\,{}_{B_{1}}\langle 1|\otimes M_{B_{2}}+|0\rangle_{B_{1}}\,{}_{B_{1}}\langle 0|\otimes I_{B_{2}},$ (71)
where $M_{B_{2}}$ is the measurement operator applied to the qubit $B_{2}$. Then, applying $W^{(4)}_{B}$ to $|\Phi_{4}\rangle$ , we select the term with the state $|1\rangle_{B_{2}}$ from this superposition, thus removing the garbage and obtaining:
$|\Phi_{5}\rangle=W^{(4)}_{B}|\Phi_{4}\rangle=|\Psi_{out}\rangle\,|0\rangle_{M_{2}}|0\rangle_{C_{1}}|0\rangle_{R_{2}}|0\rangle_{K_{2}}|1\rangle_{B_{1}}|1\rangle_{B_{2}},$ (72)
where $|\Psi_{out}\rangle$ is determined in Eq.(69). In this case, only the single run of the algorithm is enough to get the desired state $|\Psi_{out}\rangle$.
Figure 5: (a) The circuit for the matrix multiplication algorithm; the state $|\Psi_{out}\rangle$ is formed by the subsystems $R_{1}$, $C_{2}$, $K_{1}$ and $M_{1}$; $X\equiv\sigma^{(x)}$, $Z\equiv\sigma^{(z)}$; we omit subscripts in notations $W^{(j)}$ for the brevity; operator $W^{(2)}$ is a novelty in comparison with earlier algorithm in ZBQKW_arXive2024 ; (b) Notation for the multi-qubit CNOT.
V.1 Manipulations with input matrices
In this section we consider several manipulations with the input matrices $A^{(1)}$ and $A^{(2)}$. Instead of multiplication $A^{(1)}A^{(2)}$ we can perform multiplications $(A^{(1)})^{\dagger}A^{(2)}$, $A^{(1)}(A^{(2)})^{\dagger}$, $(A^{(1)})^{\dagger}(A^{(2)})^{\dagger}$, $(A^{(1)}A^{(2)})^{\dagger}=(A^{(2)})^{\dagger}(A^{(1)})^{\dagger}$ and change the order of the multipliers. With this purpose we introduce the following operators.
To organize the Hermitian conjugate of $A^{(i)}$, $i=1,2$, we use operator $W$, Eq.(46), introduced in Sec.IV.1 and rewrite it using the new notations:
@@ V.1 Manipulations with input matrices
To change the order of the matrices in the multiplication algorithm we apply the operator
$Q^{(3)}_{R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}K^{(1)}K^{(2)}}=SWAP_{R^{(1)}R^{(2)}}SWAP_{C^{(1)}C^{(2)}}SWAP_{M^{(1)}M^{(2)}}SWAP_{K^{(1)}K^{(2)}}.$ (74)
@@
Here, the multi-qubit $SWAP$-operators $SWAP_{R^{(1)}R^{(2)}}$ and $SWAP_{C^{(1)}C^{(2)}}$ establish exchange between the appropriate qubits of two selected subsystems. The circuit implementing these operators is shown in Fig.6. For instance, to calculate the product $A^{(1)}(A^{(2)})^{\dagger}$ we apply the operator $Q^{(2)}_{R^{(2)}C^{(2)}M^{(2)}}$ to the state $|\Phi_{0}\rangle$, Eq.(55). To calculate $(A^{(1)}A^{(2)})^{\dagger}$ we apply the operators $Q^{(3)}_{R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$, $Q^{(2)}_{R^{(2)}C^{(2)}M^{(2)}}$, $Q^{(1)}_{R^{(1)}C^{(1)}M^{(1)}}$ to the state $|\Phi_{0}\rangle$. Application of operators $Q^{(i)}$, $i=1,2,3$, can be controlled by additional qubits $q_{i}$ of the subsystem $q$ (see Fig.6), i.e., by the following controlled operators:
$\hat{Q}^{(i)}_{S^{(i)}}=|1\rangle_{q_{i}}\,{}_{q_{i}}\langle 1|\otimes Q^{(i)}_{S^{(i)}}+|0\rangle_{q_{i}}\,{}_{q_{i}}\langle 0|\otimes I_{S^{(i)}},$ (75)
@@
where $S^{(i)}=R^{(i)}C^{(i)}M^{(i)}$, $i=1,2$, $S^{(3)}=R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}K^{(1)}K^{(2)}$ and $I_{S^{(i)}}$ is the identity operator applied to $S^{(i)}$. The state $|\varphi\rangle_{q}=|n_{1}\rangle_{q_{1}}|n_{2}\rangle_{q_{2}}|n_{3}\rangle_{q_{3}}$, $n_{i}=0,1$ ($i=1,2,3$), of the controlling subsystem $q$ is not changed after running the algorithm.
@@
Of course calculating $(A^{(1)}A^{(2)})^{\dagger}$ can be done by performing the Hermitian conjugation after multiplication $A^{(1)}A^{(2)}$, this is another option.
@@
Figure 6: (a) Controlled manipulations with input matrices $A^{(1)}$ and $A^{(2)}$; (b) Notation for the multi-qubit SWAP-operator
V.2 Measuring normalization $G$ and discussion
@@
Although the controlled measurement allows to resolve the problem of exponential decrease of the success probability with the number of qubits encoding considered matrices, we lose the information about normalization $G$ in Eq.(69). However, this normalization might be required in obtaining the final result. Below we explore the method of probabilistic measuring the normalization $G$.
@@
Let $S_{1}=|b^{(1)}b^{(2)}|^{2}$, this value is known due to the normalization in matrix encoding, see Eq.(50). If, along with $S_{1}$, we know $S_{1}/G^{2}=\tilde{S}_{1}$, then $G=\sqrt{S_{1}/\tilde{S}_{1}}$. To find $\tilde{S}_{1}$, we refer to the state $|\Psi_{out}\rangle$, Eq.(69), and measure the state of the subsystem $K_{1}$ with the output $|0\rangle_{K_{1}}$. It is easy to understand that the probability of this output equals $\tilde{S}_{1}$, and can be found in result of multiple runs of the algorithm. It is important that the problem of small success probability is resolved on passing from $|\Phi_{4}\rangle$, Eq.(67), to $|\Phi_{5}\rangle$, Eq.(68), and does not appear in the above measurement, where the success probability $\tilde{S}_{1}$ does not depend on $n$.
@@
We have to note that the last measurement destroys the part of superposition quantum state in which the result of matrix multiplication is stored. Therefore, to get the full information, i.e., the normalization $G$ and matrix product, we have to use the set of runs to probabilistically determine $G$ and then run the algorithm one more time to end up with the state $|\Phi_{5}\rangle$ that includes $|\Psi_{out}\rangle$ encoding the product $A^{(1)}A^{(2)}$. However, if $G$ is not needed, then the single run is enough.
We emphasize that both introduced extensions do not decline the parameters of the multiplication circuit such as depth and space. Fig.5 shows that the depth is determined by the operator $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}$, while the manipulations with the input matrices in Sec.V.1, see Fig.6, increase this depth only by insignificant constant factor. Thus, the estimation of the depth remains $O(n)$ (where $N=2^{n}$) similar to the estimation in ZBQKW_arXive2024 . Regarding the space, the circuit in Fig.5 includes only five additional one-qubit subsystems ($K^{(i)}$, $M^{(i)}$, $i=1,2$ , and $B_{2}$) in comparison with the circuit in ZBQKW_arXive2024 with one of them (the ancilla qubit $B_{2}$) is needed for organization of controlled measurement. Therefore, the space estimation remains $O(n)$ as well. Of course, multiple running is required in the case of implementing the usual ancilla-measurement ZBQKW_arXive2024 . Then the depth of the whole algorithm is bigger than the depth of the circuit by $O(2^{n+1}/G^{2})$ times and equals $O(2^{n+1}n/G^{2})\sim O(2^{n}n)$.
VI Conclusion
In this paper we study some new aspects of quantum algorithms with their application to the matrix-manipulation algorithms proposed in ZQKW_2024 ; ZBQKW_arXive2024 and based on the encoding the matrix elements into the pure superposition state of a quantum system. However, applicability of introduced novelties spreads much beyond of proposed examples.
The most important issue is the replacement of the ordinary one-qubit ancilla-state measurement (needed to remove the garbage acquired during evaluation of the quantum algorithm) with the measurement of the ancilla-state controlled by the state of another one-qubit ancilla (controlled measurement). This modification allows to resolve the problem of small access probability to the desired state of the ancilla. Remember, that the problem of small success probability appears in the case of ordinary measurement of the ancilla state in HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 ; ZQKW_2024 ; ZBQKW_arXive2024 . Thus, the controlled measurement significantly increases the efficiency of the algorithms developed in those papers. At that, we lose the information on the normalization constant for the output state because this constant is expressed in terms of the access probability to the desired ancilla state and this probability is not measured in the modified algorithm. However, the normalization constant can be revealed as shown in the matrix multiplication algorithm, Sec.V.2. We also have to note that the method of practical realization of the controlled measurement is not developed yet, but its realization would be one more justification of reality of superposition quantum states.
@@ VI Conclusion
We also introduce two extensions for the quantum data encoding and apply them to the matrix-multiplication algorithm proposed in ZBQKW_arXive2024 . The first extension consists in separating the real and imaginary parts of the complex matrix under consideration by encoding these two parts into two orthogonal subspaces of the quantum state-space. Such extension allows constructing effective subroutine for operations typical for complex numbers, which are complex and Hermitian conjugations. However, this separation of real and imaginary parts requires proper modifications of the quantum algorithms, in particular, quantum algorithms for matrix addition, multiplication, determinant calculation, matrix inversion and solving linear systems that are described in ZQKW_2024 ; ZBQKW_arXive2024 . In our paper, we present such modification for the multiplication algorithm and demonstrate manipulations with the input matrices using Hermitian conjugation subroutine.
@@
Our second extension weakens the normalization constraint, imposed on the elements of the input matrices, which follows from the normalization condition for a superposition pure quantum state. This can be done by including an additional term in the superposition state encoding a particular matrix which leads to replacing the normalization equality, Eq.(43), with inequality (48) that imposes only the upper bound on the sum of the squared modulus of the matrix elements. We emphasize that the applicability of the two last extensions is not restricted to the matrix encoding, any data-set can be used instead of matrices.
@@
Of course, each of the above modifications requires involving an extra qubit. Therefore, along with four subsystems $R^{(i)}$ and $C^{(i)}$, $i=1,2$, enumerating rows and columns of the matrices $A^{(i)}$, $i=1,2$, the $i$th matrix is associated with the one-qubit subsystem $M^{(i)}$, responsible for separate encoding of the real and imaginary parts of complex matrix elements, and the one-qubit subsystem $K^{(i)}$ allowing to introduce an extra term into the superposition state encoding the matrix $A^{(i)}$, namely the presence of this extra term weakens the normalization constraint. Thus, only four additional qubits appear in the modified multiplication algorithm. In addition, involving the controlled measurement requires an additional qubit for the ancilla $B$ which becomes the two-qubit ancilla.
@@
We emphasize that all the above modifications do not significantly decline the depth and space of the circuit for multiplication algorithm. Estimations for both of them remain $O(n)$, which is similar to the appropriate characteristics of the circuit for the multiplication algorithm in ZBQKW_arXive2024 . However, the depth of the whole multiplication algorithm in ZBQKW_arXive2024 is much bigger because of the multiple running of the algorithm to access the desired ancilla-state, it equals $O(2^{n}n)$.
Proposed modifications can be included in other matrix-manipulation algorithms considered in Refs.ZQKW_2024 ; ZBQKW_arXive2024 , as well as in any measurement-based algorithm HHL ; WWLN ; ZDB ; ZBD ; Wei ; WeiBook and algorithm based on data encoding ZZRF ; BWPRWL ; A ; RML ; WBL ; SSP ; Wang2 .